Matt is 38 years younger than Carrie. 9 years ago, Carrie's age was 3 times Matt's age. How old is Matt now?

Answer by jorel1380(3719)

Let c be Carrie, and c-38 be Matt. Then
3((c-38)-9)=c-9
3c-141=c-9
2c=132
c=66
Carrie is 66; Matt is 28
☺☺☺☺